Which of the following statements are true?
The point (-2,-3) is in quadrant II of the Cartesian plane.
The equation 2x-3y=5 is a linear equation.
The distance between the points (-5,1) and (-7,9) is 2^17
^a^2 +b^2=a+b
This has me all confused, thanks for the help

Answer by praseenakos@yahoo.com(507)

QUESTION:
Which of the following statements are true?
The point (-2,-3) is in quadrant II of the Cartesian plane.
The equation 2x-3y=5 is a linear equation.
The distance between the points (-5,1) and (-7,9) is 2^17
^a^2 +b^2=a+b
------------------------------------------------------------------------------
ANSWERS:
The point ( -2, -3 ) is in quadrant II of the Cartesian plane.
No, the point ( -2, -3 ) is not in the II quadrant
Explanation:
In the I quadrant both the x and y coordinates are positive.
In the II quadrant x coordinate is negative while y coordinate is +ve.
In the III quadrant x coordinate is -ve and y coordinate is -ve.
In the IV quadrant x coordinate is +ve and y coordinate is -ve.
So the point ( -2, -3 ) lies in the III quadrant.
--------------------------------------------------------------------------------
2.The equation 2x-3y=5 is a linear equation.
Explanation:
A linear equation is an equation involving only the sum of constants or products of constants and the first power of a variable.

These equations are called "linear" because they represent straight lines in Cartesian coordinates.

A common form of a linear equation in two variables is y = mx + b,
(e.g. y = 3x + 5the line crosses the y-axis.
Equations involving terms such as x^2, y^(1/3), and (xy )are "non-linear".

The given equation 2x-3y=5 is linear because its terms involves only first power of variable.
-------------------------------------------------------------------------------
3.The distance between the points (-5,1) and (-7,9) is 2^17

Explanation:
In coordinate geometry the distance between two points is given by
Distance Formula, d = square root of {( x1 -x2)^2 + ( y1 - y2 )^2}
here

x1= -5
x2= -7
y1= 1
y2= 9
so, d = square root of {( -5 - -7)^2 + (1- 9 )^2}
= square root of {( -5+ 7)^2 + ( 1-9)^2}
= square root of {( 2)^2 + ( -8)^2}
= square root of {(4 + 64 }
= square root of { 68 }
= 2 * square root of 17
= 2 ( 17 )^1/2
-----------------------------------------------------------------------------
Last Question is not clear.



Regards,
Praseena.
praseenakos@yahoo.co.in